Sol Summary – April 17th

Crew 280 Sol Summary Report 17-04-2023

Sol: 1

Summary title: Do you copy, Earth?

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: Mariona Badenas-Agusti (Crew Commander), Carla Conejo (Executive Officer), Neus Sabaté (Mission Specialist), and Laia Ribas (Greenhab officer) woke up early in order to start preparing their scientific experiments. In the meantime, Ariadna Farrés (Health and Safety Officer), Núria Jar (Journalist), and Cesca Cufí (Engineer) drove back to the MDRS after Cesca was discharged from the hospital yesterday evening. Fortunately, she is now feeling much better and is on her way to a full recovery!

The full crew was reunited around 11:30 AM, at which point we attended our second training session. This time, we learned about EVA procedures, including how to wear the spacesuits, drive the rovers, use the radios to communicate with the Earth, and navigate the Martian terrain. At lunchtime, Mariona and Carla prepared mashed potatoes with egg as well as a broccoli & cheddar soup –a well-deserved break after a morning of intense emotions.

Crew 280 officially kicked off their Martian simulation in the early afternoon, and soon after, the team performed their first EVAs. For this purpose, they drove the rovers to Marble Ritual, where they had an opportunity to walk around, become familiar with the landscape, perform basic movements for the suit, and test their communications protocol. This evening, we are looking forward to having our first dinner on Mars with the full crew! It will be our first night on the Red Planet.

Look ahead plan: Crew 280 will receive a media visit tomorrow. We will first invite them to our group meeting, and they will then accompany us to our morning EVA (provided that Mission Support grants us permission). In the afternoon, the media team will interview members of the crew to better understand their research projects. We are convinced that it will be a busy day, but we are excited to share our mission with them so that they can help us bring space closer to society.

Anomalies in work: (i) Non-functional robotic telescope, (ii) radio #2 and #5 lost communication temporarily (we highly suspect that the headphones were slightly disconnected; our Crew Engineer and Mission Specialist have tested the radios inside the hab and they made sure that they work well — see Operations Report); (iii) the collar of helmet #2 was upside down and one metallic wire is missing (see Operations Report); (iv) the zebrafish of our Greenhab officer were killed by the high temperatures (see Greenhab Report).
Weather: Sunny, with a bit of wind in the late afternoon. Low 58F/High 75F.

Crew physical status: Our Crew Engineer is back to the station and is feeling much better.
EVA: We conducted two EVAs to Marble Ritual to put into practice all the lessons learned during our EVA training session.

Reports to be filed: Sol Summary, EVA request, EVA Report #1, EVA Report #2, HSO Report, Operations Report, Greenhab Report, Journalist Report, Pictures of the Day.

Support Requested: We would like to know how to dispose of the fish in an appropriate manner without violating state law.

Copyright © The Mars Society. All rights reserved. | Main Site